au must show your work for all problems and the answers must be in correct form SI to SI basic unit 1) 0.0327 Mm ? m 2) 742,380.56 ? m For 3 thru 11 do the conversion in either scientific form OR decimal form-not both! YOU MUST SHOW YOUR WORK to get full credit for a correct answer SI to SI 3) 5,543 mg -? kg US to US 7) 90 mph? US to Si 10) 10.0 L?gal solve the formula for the indicated variable, correct form is: indicated variable- other variables 4) 5.823 Mg = ? kg 8) 4608 in2 -? ft 11) 442 in3-?L Sl to US 5) 105 km ?ft 6) 5.00 me-?re rt 9) 30 gal - 2 ft Ho I 12) P = F for A 13) B = 20R for R 14) = Gman2 for G 15) F= 2k11 12 for 1 16) B = HolN for N 17) F.-Kaïq2 for r 16) B=HolN r2 18) F=1mU2 for U 19) A=4Lw + 2Lh + 6wh for w 2

Need help! Do I need a chart or something to help with these? thanks everyone
Answer #1

3)

5543 mg = 5543 mg (1 kg/1000000) = 0.005543 kg

4)_

5.823 Mg = 5.823 Mg (1000000 g/1 Mg) = 5823000 g

5)

105 km = 3280.84 ft

105 km = 105 km (3280.84 ft/ 1 km) = 344488.2 ft

6)

5 m3 = 5 m3 (3.28 ft/ 1 m)3 = 5 (3.28)3 ft3 = 176.4 ft3

7)

90 mph = 90 mph ((1.5 ft/s) / 1 mph ) = 135 ft/s

8)

4608 in2 = 4608 in2 (0.0833 ft/in)2 = 4608 (0.0833)2 ft2 = 31.97 ft2

9)

30 gal = 30 (0.134) ft3 = 4.02 ft3

10 )

10 L = 10 (0.264 ) gallon = 2.64 gallon